Dantzler Corporation is a fast-growing supplier of office products. Analysts project the following free cash flows (FCFs) during the next 3 years, after which FCF is expected to grow at a constant 8% rate. Dantzler's WACC is 13%.

Year 0 1 2 3
....... ....... ....... ....... ....... ....... ....... .......
....... ....... ....... ....... ....... ....... ....... ......
FCF ($ millions) - $20 $22 $39
  1. What is Dantzler's horizon, or continuing, value? (Hint: Find the value of all free cash flows beyond Year 3 discounted back to Year 3.) Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55. Do not round intermediate calculations. Round your answer to two decimal places.
    $   million
  2. What is the firm's market value today? Assume that Dantzler has zero non-operating assets. Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55. Do not round intermediate calculations. Round your answer to two decimal places.
    $   million
  3. Suppose Dantzler has $118 million of debt and 27 million shares of stock outstanding. What is your estimate of the current price per share? Write out your answer completely. For example, 0.00025 million should be entered as 250. Do not round intermediate calculations. Round your answer to the nearest cent.
    $  
Answer #1

FCF Constant Growth rate after 3 year(g) = 8%

WACC = 13%

a). Calculating Horizon Value:-

Horizon Value = \frac{FCF_3(1+g)}{(WACC-g)}

Horizon Value = \frac{39(1+0.08)}{(0.13-0.08)}

= $842.40 millions

b). Calcylating Firm's Market Value:-

Market Value FCF1 (1 + WACC) FCF2 (1 + WACC)2 FCF3 (1+WACC)3 + Horizon Value (1 +W ACC)3

Market Value = \frac{-20}{(1+0.13)^1}+\frac{22}{(1+0.13)^2}+\frac{39}{(1+0.13)^3}+\frac{842.40}{(1+0.13)^3}

Market Value = -$17.70 + $17.23 + $27.03 + $583.83

Total market value of Firm= $610.38 millions

c). Firm market value = Value of equity + Value of Debt

$610.38 milliom = Value of equity + $118 million

Value of equity = $492.38 million

Current Price per share = Value of equity/No of share outstanding

Current Price per share = $492.38 million/27 million

= $18.24 per share

If you need any clarification, you can ask in comments.   

If you like my answer, then please up-vote as it will be motivating